Grade 11Mechanics

a particle is fired vertically upward from earth’s surface and it goes up to a height equal to earth’s radius. Find the initial speed of the particle

Urvashi Kaul
10 Years agoGrade 11

1 Answer

Riddhish Bhalodia
10 Years ago
this is a simple case of energy conversation but as the heights are very high we consider the potential energy between two masses m1, m2at separation R as
E = -Gm_1m_2/R
so considering M to be mass of earth and R being it’s radius.
and m being the mass of the object thrown (with velocity v) we apply energy conservation between two points as
-GMm/R + 1/2mv^2 = -GMm/2R
1/2mv^2 = GMm/2R
v = \sqrt{GM/R}